OpenStudy (anonymous):

find f(g(x)) and g(f(x)) andetermine whether the pair of functions f and g ae inerses of each other f(x)=8x+6and g(x)=x-6/8

OpenStudy (anonymous):

they are definitely inverses, which means you should get \[f(g(x))=x\] and \[g(f(x))=x\]

OpenStudy (anonymous):

first one looks like \[f(g(x))=f(\frac{x-6}{8})=8\times (\frac{x-6}{8})+6\] \[=x-6+6=x\]